The values listed below are waiting times (in minutes) of customers at two different banks

Bank_A   Bank_B  
6.4   4.3  
6.6   5.5  
6.7   5.8  
6.8   6.2  
7.1   6.7  
7.2   7.8  
7.6   7.8  
7.8   8.4  
7.8   9.4  
7.8   10  

The values listed below are waiting times (in minutes) of customers at two different banks. At Bank A, customers enter a single waiting line that feeds three teller windows. At Bank B, customers may enter any one of three different lines that have formed at three teller windows. Answer the following questions.
Construct
aa
9999%
confidence interval for the population standard deviation
sigmaσ
at Bank A.
nothing
minless than<sigma Subscript Bank Upper AσBank Aless than< nothing
min
(Round to two decimal places as needed.)
Statistics
Answer #1

Ans:

sample standard deviation=0.543

n=10

df=10-1=9

alpha=1-0.99=0.01

alpha/2=0.005

critical chi square values:

CHIINV(0.005,9)=23.589

CHIINV(0.995,9)=1.735

99% confidence interval for the population standard deviation:

Lower limit=sqrt((10-1)*0.543^2/23.589)=0.34

Upper limit=sqrt((10-1)*0.543^2/1.735)=1.24

0.34